solve x' + (ln3)x = 0

solve x' + (ln3)x = 0
Answer #1

dt + C n x -In3t +C we know that, l0ga Cb-a -t. In 3 +Ci) we also know, ea eaeb -t. In3 Ci et, e c. k we knou, alnbs In b tInD:ferential ea (m 3) -0 dependable varable t is independent vanuble here, is So, lets assume ie n3)x In3) xo + d-t -(n3)x dif you have any doubt in solution ask me in comment section I"ll explain

Differential ear x + 1m3) a=0 here, x is deendable variable So, let's assume tis independent variable. il n=f(t) - na da :: X + (n 3) x20 - dx + (113)x=0 5 dl = -(\n3)x dx = 4n3). dt at integrate both sides. 5 s due = 103.60t +C. _(s is constant) Ina=-11n3.t +ci x= 113.(-+)+(1) n=edingt = x=3 tec -) [a= 3? c. cis arbitarasy constant), (c= e.